# Velocity rotation curves in

### v1

@@ 1 Introduction
General relativity admits a wide range of exact solutions, many of which display non-trivial structures even in the absence of matter. Among these, the gravimagnetic dipole spacetime, a recently described [1, 2, 3, 4] axisymmetric, asymptotically flat and stationary metric, consists of two counter-rotating NUT (Newman-Unti-Tamburino) black holes, of equal mass but opposite NUT charges, and connected by a Misner string. When the separation between the black holes is tuned precisely, the string becomes tensionless and the configuration is at equilibrium. By adimensionalising the system, the mass scale is removed and the gravimagnetic dipole spacetime is characterised by only one parameter, the NUT charge.
In this work, the motion of massive and massless test particles is studied in the equatorial plane, making full use of the available symmetries. Using a Hamiltonian formalism, an effective potential[5] is defined, leading to a condition for the existence of circular orbits. The rotation velocities associated with these orbits are then computed, thereby obtaining the velocity rotation curves for various values of the NUT parameter.
The recent study of the gravito-electromagnetic approximation to the gravimagnetic dipole [6] derived an approximate velocity rotation curve for circular orbits within the weak-field regime of the gravimagnetic dipole spacetime, identifying conditions under which a roughly flat rotation curve emerges without invoking dark matter. The present exact results are compared with these approximate analytic expressions, in the relevant domain of parameters.
@@
Section 4 finally presents the calculation of the velocity rotation curves along with the corresponding figures. Comparison is made with the results from [6], showing good agreement in the relevant domain of parameters.
@@ 2 The gravimagnetic dipole spacetime
Let us consider the gravimagnetic dipole spacetime metric. This configuration [1, 4] consists of the nonlinear superposition of two counter-rotating NUT objects of equal masses $m>0$ and opposite NUT charges $\pm\nu$ ($\nu\geq 0$), separated by a total distance $2k\geq 2m>0$, and positioned symmetrically on the $z$-axis relative to $z=0$.
@@ 2.1 The values of $m$, $k$ and $\nu$
Figure 2: Cut of the spacetime along the $XZ$ plane for $\nu=0.1$. The orange dots represent the black holes. The blue lines represent the static limit. The dashed grey lines represent the horizon rods.
Figure 3: Cut of the spacetime along the $XZ$ plane for $\nu=0.8$. The orange dots represent the black holes. The blue lines represent the static limit. The dashed grey line represent the horizon rods.
@@
On the condition that $\alpha_{\pm}$ and $d$ are real, the system defined by the gravimagnetic dipole spacetime consists of two black holes separated by a distance $2k$ and connected by a spinning cosmic string – called Misner string – of length $2\alpha_{-}$. These correspond to three sections of the $z$ axis; the two black holes horizons for $\alpha_{-}<z<\alpha_{+}$ and $\alpha_{+}<z<-\alpha_{-}$, and the Misner string for $\alpha_{-}<z<\alpha_{-}$. The masses of the black holes $M_{H_{\pm}}$ and their angular momenta, $J_{H_{\pm}}$ are functions of the three parameters $m,k$ and $\nu$ [4].
The conditions for $\alpha_{\pm}$ and $d$ to be real and non zero are given[4] by
@@
Lastly, for $k=m$, it was shown[1] that the system describes a Kerr black hole of mass $2m$ and Kerr parameter $\nu=a$, provided that $\nu<2m$. In this case, $d=mk$, $\alpha_{+}$ is real and $\alpha_{-}$ is imaginary, but $\alpha_{-}$ will not appear in any of the expressions of the metric, preserving its reality.
@@ 3 The effective potential for geodesics
As is well known[7, 8, 9], the particle’s dynamics is derived from the Hamiltonian first-order action:
@@ 3.3 Massive particules
Figure 5: For different values of the energy $E$, as functions of the parameter $\nu$, the radii of the possible circular orbits $\rho_{\pm}$ for massive particles. Depending on the value of the NUT parameter $\nu$ and the energy $E$, there is from 0 to $4$ possible orbits.
@@ 3.4 Massless particles
Figure 6: The top (resp. bottom) graph represents the first (resp. second) derivative of the potential as a function of $\rho$ for massless particles, for two different rotation directions (prograde and retrograde) and two different values of $\nu$. The number of possible circular orbits (from 0 to $4$) depends on this parameter.
@@ 4 Proper distance and velocity
The radar distance, or spatial distance[11] is the distance determined by the following procedure: an observer at point $B$ sends a light signal to point $A$ and receives its reflection back after some short interval of time. The radar distance between $A$ and $B$ is defined as half the total travel time (measured in the observer’s proper time) multiplied by the speed of light.
@@ 4.3 The massive case
In Fig. 11 the velocity rotation curve, computed at $\nu=0.999$, is compared with the approximate expression given in eq. (79) of [6]. In this regime ($\nu\approx 1$, tensionless Misner string), where the gravito-electromagnetic approximation is expected to hold, we find close agreement between the two results. (For $\nu=0.999$ the following values are found: $k=44.716$, $\alpha_{+}=46.107,\alpha_{-}=43.279$.)
Figure 11: For a value of the NUT parameter of $\nu=0.999$, the blue plain (resp. dashed) line depicts the exact velocity of a massive particle on a prograde (resp. retrograde) circular orbit. The red dashdotted line depicts the gravito-electromagnetic approximation in eq. (79) of Ref. [6].

### v2

@@ 1 Introduction
General relativity admits a wide range of exact solutions, many of which display non-trivial structures even in the absence of matter. Among these, the gravimagnetic dipole spacetime, a recently described [undef, undefa, undefb, undefc] axisymmetric, asymptotically flat and stationary metric, consists of two counter-rotating NUT (Newman-Unti-Tamburino) black holes, of equal mass but opposite NUT charges, and connected by a Misner string. When the separation between the black holes is tuned precisely, the string becomes tensionless and the configuration is at equilibrium. By adimensionalising the system, the mass scale is removed and the gravimagnetic dipole spacetime is characterised by only one parameter, the NUT charge.
In this work, the motion of massive and massless test particles is studied in the equatorial plane, making full use of the available symmetries. Using a Hamiltonian formalism, an effective potential[undefd] is defined, leading to a condition for the existence of circular orbits. The rotation velocities associated with these orbits are then computed, thereby obtaining the velocity rotation curves for various values of the NUT parameter.
The recent study of the gravito-electromagnetic approximation to the gravimagnetic dipole [undefe] derived an approximate velocity rotation curve for circular orbits within the weak-field regime of the gravimagnetic dipole spacetime, identifying conditions under which a roughly flat rotation curve emerges without invoking dark matter. The present exact results are compared with these approximate analytic expressions, in the relevant domain of parameters.
@@
Section 4 finally presents the calculation of the velocity rotation curves along with the corresponding figures. Comparison is made with the results from [undefe], showing good agreement in the relevant domain of parameters.
@@ 2 The gravimagnetic dipole spacetime
Let us consider the gravimagnetic dipole spacetime metric. This configuration [undef, undefc] consists of the nonlinear superposition of two counter-rotating NUT objects of equal masses $m>0$ and opposite NUT charges $\pm\nu$ ($\nu\geq 0$), separated by a total distance $2k\geq 2m>0$, and positioned symmetrically on the $z$-axis relative to $z=0$.
@@ 2.1 The values of $m$, $k$ and $\nu$
Figure 2: Cut of the spacetime along the XZXZ plane for $\nu=0.1$. The orange dots represent the black holes. The blue lines represent the static limit. The dashed grey lines represent the horizon rods.
Figure 3: Cut of the spacetime along the XZXZ plane for $\nu=0.8$. The orange dots represent the black holes. The blue lines represent the static limit. The dashed grey line represent the horizon rods.
@@
On the condition that $\alpha_{\pm}$ and $d$ are real, the system defined by the gravimagnetic dipole spacetime consists of two black holes separated by a distance $2k$ and connected by a spinning cosmic string – called Misner string – of length $2\alpha_{-}$. These correspond to three sections of the $z$ axis; the two black holes horizons for $\alpha_{-}<z<\alpha_{+}$ and $\alpha_{+}<z<-\alpha_{-}$, and the Misner string for $\alpha_{-}<z<\alpha_{-}$. The masses of the black holes $M_{H_{\pm}}$ and their angular momenta, $J_{H_{\pm}}$ are functions of the three parameters $m,k$ and $\nu$ [undefc].
The conditions for $\alpha_{\pm}$ and $d$ to be real and non zero are given[undefc] by
@@
Lastly, for $k=m$, it was shown[undef] that the system describes a Kerr black hole of mass $2m$ and Kerr parameter $\nu=a$, provided that $\nu<2m$. In this case, $d=mk$, $\alpha_{+}$ is real and $\alpha_{-}$ is imaginary, but $\alpha_{-}$ will not appear in any of the expressions of the metric, preserving its reality.
@@ 3 The effective potential for geodesics
As is well known[undeff, undefg, undefh], the particle’s dynamics is derived from the Hamiltonian first-order action:
@@ 3.3 Massive particules
Figure 5: For different values of the energy $E$, as functions of the parameter $\nu$, the radii of the possible circular orbits $\rho_{\pm}$ for massive particles. Depending on the value of the NUT parameter $\nu$ and the energy $E$, there is from 0 to 44 possible orbits.
@@ 3.4 Massless particles
Figure 6: The top (resp. bottom) graph represents the first (resp. second) derivative of the potential as a function of $\rho$ for massless particles, for two different rotation directions (prograde and retrograde) and two different values of $\nu$. The number of possible circular orbits (from 0 to 44) depends on this parameter.
@@ 4 Proper distance and velocity
The radar distance, or spatial distance[undefj] is the distance determined by the following procedure: an observer at point $B$ sends a light signal to point $A$ and receives its reflection back after some short interval of time. The radar distance between $A$ and $B$ is defined as half the total travel time (measured in the observer’s proper time) multiplied by the speed of light.
@@ 4.3 The massive case
In Fig. 11 the velocity rotation curve, computed at $\nu=0.999$, is compared with the approximate expression given in eq. (79) of [undefe]. In this regime ($\nu\approx 1$, tensionless Misner string), where the gravito-electromagnetic approximation is expected to hold, we find close agreement between the two results. (For $\nu=0.999$ the following values are found: $k=44.716$, $\alpha_{+}=46.107,\alpha_{-}=43.279$.)
Figure 11: For a value of the NUT parameter of $\nu=0.999$, the blue plain (resp. dashed) line depicts the exact velocity of a massive particle on a prograde (resp. retrograde) circular orbit. The red dashdotted line depicts the gravito-electromagnetic approximation in eq. (79) of Ref. [undefe].